OpenStudy (anonymous):

could someone help with this question a hockey puck hits the boards at a velocity of 12m/s at an angle of 30 deg in the boards, it is deflected with a velocity of 10m/s at an angle of 25 deg to the boards. determine the pucks change in velocity

OpenStudy (anonymous):

is anyone able to start me off

OpenStudy (anonymous):

the above goes with the following question an airplane flies with a headin of [N58degW] a distance of 618km. then it flies 361km deading [E35degS]. determine the planes displacement of trip

OpenStudy (anonymous):

this is what I got so far dax = 618(cos58)= 327 day = 618(sin58)= 524 dbx = 361(cos35)= 296 dby=361(sin35)207= 207 Rx = dax + dbx = 623 Ry = day + dby = 731 R^2 = (Rx)^2 + (Ry)^2 = 922496 R= 960 is this right so far